OpenStudy (firejay5):

The formula for the perimeter of a rectangle is P = 2w + 2l. If l = 2 in. and w = 4 in., what operation(s) would you do first?

OpenStudy (goformit100):

first take 2 out as common. like this :- P = 2w + 2l P = 2x4 + 2x2 P = 2(4+2) P = 2(6) = 2x6 = 12 in.